OpenStudy (anonymous):

how do you write this in standard form 8, -14, and 3 + 9i

OpenStudy (anonymous):

what do you mean by standard form?

OpenStudy (anonymous):

you do know that you can't have a function with these roots right?

OpenStudy (anonymous):

I have to turn those roots into something like this f(x) = x4 - 11x3 + 72x2 - 606x + 10,080

OpenStudy (anonymous):

the number of complex roots of a function is always even. That's why it is difficult to deal with functions higher than degree 5.

OpenStudy (anonymous):

given that you have forgotten to provide me the root 3-9i, your answer would be: f(x)=(x-8)(x+14)(x-3-9i)(x-3+9i) If you are still insisting on not having the root 3-9i, your answer will be the first three parentheses.

OpenStudy (anonymous):

clear on that?

OpenStudy (anonymous):

I'm suppose to write a polynomial function of minimum degree with real coeffecients whose zeros include the numbers I gave in standard form.

OpenStudy (anonymous):

I told ya, you can't have function with odd number of imaginary roots. That's why you will be getting the imaginary element in your answer.

OpenStudy (anonymous):

Oh and btw 3-9i wasn't one of the roots given to me on my question which was why I didn't provide you with it.

OpenStudy (anonymous):

So none of these could possibly be the answer? f(x) = x4 - 11x3 + 72x2 - 606x + 10,080 f(x) = x4 - 303x2 + 1212x - 10,080 f(x) = x4 - 11x3 - 72x2 + 606x - 10,080 f(x) = x4 - 58x2 + 1212x - 10,080

OpenStudy (anonymous):

look, you have a function of degree 4, which means that you must have four roots. you provided me with only three roots and I told you that the only way that it would make sense is that you have the final root that I provided you with. without that root, you are stuck.

OpenStudy (anonymous):

If you want to check the validity of my answer, multiply the parentheses and see which one you would get.

OpenStudy (anonymous):

of course, you should keep the fact in mind that in a quiz, midterm, homework, etc, you won't be given the full thing; they should keep something in their sleeves to test you with...

OpenStudy (anonymous):

This missing root was your missing piece of puzzle.

OpenStudy (anonymous):

multiply the parentheses and tell me what you get

OpenStudy (anonymous):

okay hang on

OpenStudy (anonymous):

I got x^4 - 11x^3 - 72x^2 + 606x - 10,080

OpenStudy (anonymous):

which is one of your options, isn't it?

OpenStudy (anonymous):

Yes, am I correct?

OpenStudy (anonymous):

You are, as well as me. You did multiply all FOUR roots, right? the three you had and the one I provided you with?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then that is your answer...